OpenStudy (babynini):

Exact solutions in interval [0, 2pi)

OpenStudy (babynini):

Find the exact solutions of the equation in the interval [0, 2pi) sin 3(theta) - sin 6(theta) = 0

OpenStudy (babynini):

@rvc do you mind if I bug you with one more question? :)

OpenStudy (babynini):

(Sorry, I know everyone's asking you right now xP)

OpenStudy (babynini):

....

OpenStudy (aaronandyson):

"Just came to check up on you, and encourage you. It's late so everyone's pretty much off of openstudy and here we are. The few faithful, just trying to get our work done xD Whether your're procrastinating (like I did...not a good idea) or being extremely diligent and working late. I don't know. But keep up the hard work!! You're doing awesome. It may get harder from here, or maybe easier, idk. But it will make you a better person in the end. Take time to enjoy the sunshine and smell the flowers :)" Hehe xD

OpenStudy (aaronandyson):

And,I don't know the solution of your Question either.

rvc (rvc):

im sorry @Babynini im sure @perl will help u im busy right now :(

OpenStudy (perl):

did you try to expand $$\Large \sin(6\theta) = \sin(2\cdot 3\theta) = 2 \sin(3\theta)\cos(3\theta) $$

OpenStudy (perl):

$$ \Large { \sin (3\theta) - \sin (6\theta) = 0 \\\sin (3\theta) - 2 \sin(3\theta)\cos(3\theta)= 0 \\\sin (3\theta) (1 - 2 \cos(3\theta))= 0 }$$

OpenStudy (perl):

can you solve it now?

OpenStudy (welshfella):

recall the concept if a and b are functions and ab = 0 then a = 0 or b = 0

OpenStudy (babynini):

Em..no i'm not quite sure how to solve from there :/

OpenStudy (babynini):

@rvc :P (sorry, if you're busy again.)